OpenStudy (anonymous):

I NEED HELP!!!!!! a ball is thrown into the air with an upward velocity of 20 feet per second. Its height, h, in feet after t seconds is given by the function h(t)= -16t^2+20t+2. How long does it take the ball to reach its maximum height? What is the balls maximum height? Round to the nearest hundredth if necessary.

OpenStudy (anonymous):

Try the following equations of motion under vertical projection; v= u-gt where if you manipulate it assuming v=0 you will get t=u/g as your time of flight and as for Hmax use equation v^2=u^2-2gHmax and make Hmax the subject of the formula knowing at the back of our minds that @Hmax the value of v^2=0.

OpenStudy (anonymous):

ohh ok thank you

OpenStudy (anonymous):

max at t = -b/2a = -20/-32 = 0.63 sec Max height = h = -16(0.63)^2 + 20(6.3) _ 2 = ...